OpenStudy (anonymous):

Minimize the function f(x,y)=x²+y²-xy subject to the constraint x+2y-14=0 (we are studying the Lagrangian Multiplier if that helps!) Minimize the function f(x,y)=x²+y²-xy subject to the constraint x+2y-14=0 (we are studying the Lagrangian Multiplier if that helps!) @Mathematics

OpenStudy (amistre64):

add in an L then

OpenStudy (amistre64):

f(x,y,L) =x²+y²-xy+L(x+2y-14) and take all your partial derivatives set to 0

OpenStudy (amistre64):

fx = 2x -y+L = 0 fy = 2y - x + 2L = 0 fL = x+2y-14 = 0 but replace x and y by terms of L

OpenStudy (anonymous):

what do you mean by replace x and y by terms of L?

OpenStudy (amistre64):

fx = 2x -y+L = 0 2x = y-L x = (y-L)/2 fy = 2y - x + 2L = 0 x = 2y+2L 2y+2L = (y-L)/2 4y +4L = y-L 4y-y = -L-4L 3y = -5L y = -5L/3 x = 2(-5L/3)+2L x = -10L/3 +2L x = -10L+6L/3 x = -4L/3

OpenStudy (amistre64):

fL = x+2y-14 = 0 fL = (-4L/3)+2(-5L/3)-14 = 0 fL = -4L/3 -10L/3 -14 = 0 fL = -14L/3 -14 = 0 fL = -14L/3 = 14 fL = -14L = 42 fL = L = 42/-14 = -3

OpenStudy (amistre64):

any of that make sense?

OpenStudy (anonymous):

yes okay so I sub x and y into the parrtial derivative of L?

OpenStudy (amistre64):

essentially you solve a system of equations between fx and fy so that they are in terms of L, your lagrange multiplier then use those in your constraint to determine the value for your L

OpenStudy (amistre64):

then back substitute it to determine x and y

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

if i did it right, x = 4 and y = 5; otherwise i missed a sign and they might need to go negative